OpenStudy (anonymous):

select the equation that is perpendicular to y=-1/4x+11 and passes through point (3,5)

OpenStudy (texaschic101):

this looks familiar...have you done this question before ?

OpenStudy (anonymous):

I realized that I put in the slope wrong, and it was supposed to be negative so kind of

OpenStudy (texaschic101):

ok....the slope of the equation is -1/4. For perpendicular line we need the negative reciprocal which is 4 (" flip " the slope and change the sign) y = mx + b (solve for b) slope(m) = 4 (3,5) x = 3 and y = 5 now sub y = mx + b 5 = 4(3) + b 5 = 12 + b 5 - 12 = b -7 = b perpendicular line is : y = 4x - 7

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no. Thank you

OpenStudy (texaschic101):

sure..glad to help :)